# Discrete Fourier Transform versus Discrete Chi-Square Method

### v1

@@ Abstract
We compare two time series analysis methods, the Discrete Fourier Transform (DFT) and our Discrete Chi-square Method (DCM). DCM is designed for detecting many signals superimposed on an unknown trend. The solution for the non-linear DCM model is an ill-posed problem. The backbone of DCM is the Gauss-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models. DCM is a simple numerical time series analysis method that performs a massive number of linear least squares fits. We show that our numerical solution for the DCM model fulfills the three conditions of a well-posed problem: existence, uniqueness and stability. The Fisher-test is used to identify the best DCM model from all alternative tested DCM models. The correct DCM model must also pass our Predictivity-test. Our analyses of seven different simulated data samples expose the weaknesses of DFT and the efficiency of DCM. The DCM signal and trend detection depend only on the sample size and the accuracy of data. DCM is an ideal forecasting method because the time span of observations is irrelevant. We recommend fast sampling of large high quality datasets and the analysis of those datasets using numerical DCM parallel computation Python code.
@@ 1 Introduction
Stability means that the solution behaves predictably as the input varies. In other words, there exists a continuous mapping from the input space to the solution space.
If any of these conditions are not satisfied, the problem is classified as ill-posed. Ill-posed problems are encountered in various fields of science [22]. For example, Piskunov et al. [24] solves the stellar surface imaging inverse problem using the [29] regularization technique. Ill-posed problems are also encountered in time series analysis of astronomical data [30, 1]. The problem is ill-posed because the unknown signal frequencies are free parameters of non-linear time series analysis models. The Lomb-Scargle periodogram is the most frequently applied time series analysis method in astronomy because the observations are often unevenly spaced in time [21, 27]. Several extentions of this method have been published [35, 32]. The DFT pre-whitening technique can detect many signals from unevenly spaced data [e.g., 26, 34]. The data must be detrended before applying this technique, which detects one frequency at a time, until it detects no new frequency. In this study, artificial simulated data having equal weights is analysed using the DFT version published by Horne and Baliunas [11]. We compare the performance of this DFT version to the performance of our DCM.
[19] formulated the least squares method. [7] connected this method to the principles of probability and the normal distribution. A few years later, he showed that the least-squares method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [8]. The extended Gauss-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [33]. This Gauss-Markov theorem is the backbone of our DCM.
@@ 2 Methods
The observations are $y_{i}=y(t_{i})\pm\sigma(t_{i})$, where $t_{i}$ are the observing times and $\sigma_{i}$ are the errors (i=1,2,,n)(i=1,2,...,n). The time span of data is $\Delta T=t_{n}-t_{1}$. The mid point is $t_{\mathrm{mid}}=t_{1}+\Delta T/2$. The mean and the standard deviation of all $y_{i}$ values are denoted with $m$ and $s$.
@@ 2.1 Discrete Chi-square method (DCM)
where the integer values $K_{1}$, $K_{2}$ and $K_{3}$ are called the model orders [14]. The notation $g_{K_{1},K_{2},K_{3}}(t)$ is used to specify these orders. The DCM model is a sum of two functions. These functions are the periodic function
@@
where $n_{f}=n_{L}$ and $n_{f}=n_{S}$, respectively.
@@
The short search gives the best $f_{\mathrm{1,best}},$..., $f_{\mathrm{K_{1},best}}$ frequencies for the data. These frequencies are the unique initial values for the first group of free parameters $\bm{\beta}_{\mathrm{I,initial}}=$ $f_{\mathrm{1,best}},$..., $f_{\mathrm{K_{1},best}}]$ (Equation 9). The linear model for these constant $f_{\mathrm{1,best}},...,f_{\mathrm{K_{1},best}}]$ frequencies gives the unique initial values for the second group $\bm{\beta}_{\mathrm{II,initial}}$ of free parameters (Equation 10). The non-linear iteration
@@
  • $M_{k}$ = Polynomial coeffients.
We use the bootstrap technique to estimate the DCM model parameter errors [5]. The tested bootstrap frequencies are the same as in the short search. Our TSPA- and CPS-methods applied the same bootstrap tehnique [15, 20]. We select a random sample $\bm{\epsilon}^{*}$ from the residuals $\bm{\epsilon}$ of the DCM model (Equation 12). Any $\epsilon_{i}$ value can be chosen to the $\bm{\epsilon}^{*}$ sample as many times as the random selection happens to favour it. We create numerous artificial bootstrap data samples
@@
There are, of course, totally wrong DCM models for the data. For example, DCM can be forced search for too few or too many $K_{1}$ signals, or the selected $p(t)$ trend order $K_{3}$ can be wrong, as shown in Figures 5-10 by [13]. Such DCM models are unstable and we denote them with $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”, like in [14] and Jetsu [^1] Jetsu L. “Do planets cause the sunspot cycle?”, submitted to Scientific Reports. August 13th, 2025. (2025). These unstable models have three signatures
@@
Figure 1: Model 1 (Table 1: $n=50$, ${\mathrm{SN}}=50$ simulation). (a) DCM long search periodogram $z_{1}(f_{1})$ gives best period at 1.843 (diamond). (b) DCM short search periodogram $z_{1}(f_{1})$ gives best period at 1.822 (diamond). (c) DCM model $g(t)$ (black continuous line), DCM trend $p(t)$ (black dashed line) and data $y_{i}$ (black dots). (d) DCM model detrended g(t)p(t)g(t)-p(t) (black continuous line), DCM signal $h_{1}(t)$ (red thick continuous line), detrended data $y(t_{i})-p(t_{i})$ (black dots) and DCM model residuals $y(t_{i})-g(t_{1})$ (blue dots) offset to -0.65 level (blue dotted line). (e) DFT periodogram $z_{\mathrm{DFT}}(f)$ gives best period at 1.190 (Diamond). (f) DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line), DFT trend $p_{\mathrm{DFT}}(t)$ (black dashed line) and data $y_{i}$ (black dots). (g) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black contiuous line), DFT pure sine $s_{\mathrm{DFT}}(t)$ (red thick continuous line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -1.5 level (blue dotted line).
@@ 2.2 Simulated DCM model data samples
We use seven different $g(t)$ models to simulate 21 artificial data samples (Sections 3.1-3.7). The time interval of all simulated samples is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$. The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$. Hence, the distance between independent frequencies [18, 27, 25] is always
@@
2.3 Dicrete Fourier Transform (DFT)
@@
The data sample is “too short” [27, 28]
@@
The data contain “a trend” [3, 17]
@@
The signal frequencies are “too close” [18, 23].
@@
The signals are not “pure sines” [16, 2].
@@
We search for many pure sine signals using the DFT pre-whitening technique [26, 34]. The $p_{i}=p_{\mathrm{DFT}}(t_{i})$ trend is first removed from the data $y_{i}=y(t_{i})$. In the search for the first signal, we apply the DFT to the detrended $y_{i}-p_{i}$ data. When searching for the second signal, we apply the DTF to the original model residuals $\epsilon_{i}=y_{i}-g_{\mathrm{y,DFT}}(t_{i})$.
@@ 3.4 Model 4
Figure 7: Model 7 (Table 7: $n=100$, ${\mathrm{SN}}=5~000~000$ simulation). otations as in Figure 3. Best periods (diamonds) are explained in Section 3.7.
@@ 3.6 Model 6
We conclude that the DFT “detects” the $P_{1}/2$ and $P_{1}$ periods, where $P_{1}$ is the correct simulated period value. However, the DFT two signal model is not the correct model for these Model 6 simulated data, which contain only one signal. If the correct period is $P$ and the correct model is a double wave $(K_{2}=2)$, the DCM pure sine model $(K_{1}=1)$ analysis can also give the values $P/2$ and $P$ [14].
@@ 3.7 Model 7
Table 8: Fisher tests. DCM analysis of Model 1 simulated data ($n=100,{\mathrm{SN}}=100$, data file Model1n100SN100). (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name. (2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.
@@
22footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
@@ 3.8 Best model
Table 9: Fisher tests for Model 1 $n=100,{\mathrm{SN}}=100$ simulated data (electronic data file Model1n100SN100). (1) Polynomial model. (2) DCM model $g_{1,1,0}$. Note that $\mathcal{M}$=1-3  polynomials represent simple models, and $\mathcal{M}$=5-8  polynomials represent complex models. Otherwise as in Table 8.
@@
33footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.
In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$. We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large. One example of testing the $K_{2}=1$ and $K_{2}=2$ signal models against each other can be found in Jetsu [14, Table 4].
@@ 3.9 Significance estimates
There is no need to discuss the DFT significance estimates [11, their Equation 22] because this method fails to detect the correct Model 1 period.
@@ 3.10 Prediction
There are numerous techniques for predicting (forecasting) a time series [e.g., 10, 12]. The DCM model $g(t_{i},\bm{\beta})$ can be used to predict. We divide all data into the predictive data and the predicted data. The time points, the observations and the errors of these samples are
@@
during the chosen time interval $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$. Jetsu (2025) used this $m_{\mathrm{pred}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720 [31]. Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
@@ 4 Discussion
The solution for the non-linear DCM model is an ill-posed problem (Equations 1-5). We present a numerical solution that fulfills the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem. The DCM model is just one special case of a non-linear model. Our technique can be applied to solve other non-linear models: Use the free parameters that make the model non-linear (Equation 9: $\bm{\beta}_{I}$) for solving the remaining other free parameters (Equation 10: $\bm{\beta}_{II}$).
@@ 4.1 Existence ( $\mathrm{C}_{1}$)
[6] transformed the original function into the frequency domain. The modern DFT time series analysis method transforms the original data into the frequency domain and gives the best frequency for a pure sine model. [8] presented the least squares method, which minimizes the differences between the data and the linear model. Our DCM does the same by testing a large number of linear models. For every chosen DCM model, the number of tested linear models is

### v2

@@ Abstract
We compare two time series analysis methods, the Discrete Fourier Transform (DFT) and our Discrete Chi-square Method (DCM). DCM is designed for detecting many signals superimposed on an unknown trend. The solution for the non-linear DCM model is an ill-posed problem. The backbone of DCM is the Gauß-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models. DCM is a simple numerical time series analysis method that performs a massive number of linear least squares fits. Hence, the data spacing, even or uneven, is irrelevant. We show that our numerical solution for the DCM model fulfils the three conditions of a well-posed problem: existence, uniqueness and stability. The Fisher-test is used to identify the best DCM model from all alternative tested DCM models. The correct DCM model must also pass our Predictivity-test. Our analyses of seven different simulated data samples expose the weaknesses of DFT and the efficiency of DCM. The DCM signal and trend detection depend only on the sample size and the accuracy of data. DCM is an ideal forecasting method because the time span of observations is irrelevant. If the Gauß-Markov theorem is valid, DCM can not fail. We recommend fast sampling of large high quality datasets and the analysis of those datasets using numerical DCM parallel computation Python code.
@@ 1 Introduction
[21] formulated the least squares method. [11] connected this method to the principles of probability and the normal distribution. A few years later, he showed that the least-squares method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [12]. The extended Gauß-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [30]. The Gauß-Markov theorem is the backbone of our Discrete Chi-Square Method (DCM). This time series analysis method performs a massive number of linear least squares fits to find the best model for the data.
@@
Stability means that the solution behaves predictably as the input varies. In other words, there exists a continuous mapping from the input space to the solution space. If any of these conditions are not satisfied, the problem is classified as ill-posed.
[…]
However, no published solution seems to exist for the more complicated cases, where the signal periods are longer than $\Delta T$ and/or these signals are not pure sines. We will show that DCM can find a well-posed numerical solution even for such complicated cases. This method can detect an unknown number of periodic signals superimposed on an unknown aperiodic trend. The noise level can be unknown because the performance of DCM depends only on the sample size $(n)$ and data accuracy $(\sigma)$. The time span $(\Delta T)$ is irrelevant. The data spacing, even or uneven, is also irrelevant.
In this study, we apply one published DFT version [17] to analyse artificial simulated unevenly spaced data having equal weights. The performance of this DFT version is compared to the performance of our DCM.
@@ 2 Methods
The observations are $y_{i}=y(t_{i})\pm\sigma(t_{i})$, where $t_{i}$ are the observing times and $\sigma_{i}$ are the errors (i=1,2,,n)(i=1,2,...,n). The time span of data, which is also called the data window, is $\Delta T=t_{n}-t_{1}$. The mid point is $t_{\mathrm{mid}}=t_{1}+\Delta T/2$. The mean and the standard deviation of all $y_{i}$ values are denoted with $m$ and $s$.
@@ 2.1 Discrete Chi-square method (DCM)
where the integer values $K_{1}$, $K_{2}$ and $K_{3}$ are called the model orders. The notation $g_{K_{1},K_{2},K_{3}}(t)$ is used to specify these orders. The DCM model is a sum of two functions. These functions are the periodic function
@@
where $n_{f}=n_{L}$ and $n_{f}=n_{S}$, respectively. The even or uneven data spacing is irrelevant because the linear least squares fit result for every tested frequency combination does not depend on this spacing.
@@
The short search gives the best $f_{\mathrm{1,best}},$..., $f_{\mathrm{K_{1},best}}$ frequencies for the data. These frequencies are the unique initial values for the first group of free parameters $\bm{\beta}_{\mathrm{I,initial}}=$ $f_{\mathrm{1,best}},$..., $f_{\mathrm{K_{1},best}}]$ (Equation 9). The linear model for these constant $f_{\mathrm{1,best}},...,f_{\mathrm{K_{1},best}}]$ frequencies gives the unique initial values for the second group $\bm{\beta}_{\mathrm{II,initial}}$ of free parameters (Equation 2.1). The non-linear iteration
@@
  • $M_{k}$ = Polynomial coefficients.
We use the bootstrap technique to estimate the DCM model parameter errors [9]. The tested bootstrap frequencies are the same as in the short search. We select a random sample $\bm{\epsilon}^{*}$ from the residuals $\bm{\epsilon}$ of the DCM model (Equation 12). Any $\epsilon_{i}$ value can be chosen to the $\bm{\epsilon}^{*}$ sample as many times as the random selection happens to favour it. We create numerous artificial bootstrap data samples
@@
There are, of course, totally wrong DCM models for the data. For example, DCM can be forced search for too few or too many $K_{1}$ signals, or the selected $p(t)$ trend order $K_{3}$ can be wrong, as shown in Figures 5-10 by [19]. Such DCM models are unstable and we denote them with $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”, like in Jetsu [^1] Jetsu L. “Do planets cause the sunspot cycle?”, submitted to Scientific Reports. August 13th, 2025. (2025). These unstable models have three signatures
@@
Figure 1: Model 1 (Table 1: $n=50$, ${\mathrm{SN}}=50$ simulation). (a) DCM long search periodogram $z_{1}(f_{1})$ gives best period at 1.843 (diamond). (b) DCM short search periodogram $z_{1}(f_{1})$ gives best period at 1.822 (diamond). (c) DCM model $g(t)$ (black continuous line), DCM trend $p(t)$ (black dashed line) and data $y_{i}$ (black dots). (d) DCM model detrended g(t)p(t)g(t)-p(t) (black continuous line), DCM signal $h_{1}(t)$ (red thick continuous line), detrended data $y(t_{i})-p(t_{i})$ (black dots) and DCM model residuals $y(t_{i})-g(t_{1})$ (blue dots) offset to -0.65 level (blue dotted line). (e) DFT periodogram $z_{\mathrm{DFT}}(f)$ gives best period at 1.190 (Diamond). (f) DFT model $g_{\mathrm{DFT}}(t)$ (black continuous line), DFT trend $p_{\mathrm{DFT}}(t)$ (black dashed line) and data $y_{i}$ (black dots). (g) DFT model detrended $g_{\mathrm{DFT}}(t)-p_{\mathrm{DFT}}(t)$ (black continuous line), DFT pure sine $s_{\mathrm{DFT}}(t)$ (red thick continuous line), detrended data $y(t_{i})-p_{\mathrm{DFT}}(t_{i})$ (black dots) and DFT model residuals (blue dots) offset to -1.5 level (blue dotted line).
@@ 2.2 Simulated DCM model data samples
We use seven different $g(t)$ models to simulate 21 artificial data samples (Sections 3.1-3.7). The time interval of all simulated samples is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$. The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$. Hence, the distance between independent frequencies [20] is always
@@
2.3 Discrete Fourier Transform (DFT)
@@
The data sample is “too short” [20]
@@
The data contain “a trend” [23]
@@
The signal frequencies are “too close” [20].
@@
The signals are not “pure sines” [5].
@@
The $p_{i}=p_{\mathrm{DFT}}(t_{i})$ trend is first removed from the data $y_{i}=y(t_{i})$ [23]. Then, the technique described by Ghaderpour et al. [13] is used to detect the pure sine signals. We apply the DFT to the detrended $y_{i}-p_{i}$ data. When searching for the second signal, we apply the DTF to the original model residuals $\epsilon_{i}=y_{i}-g_{\mathrm{y,DFT}}(t_{i})$.
@@ 3.4 Model 4
Figure 7: Model 7 (Table 7: $n=100$, ${\mathrm{SN}}=5~000~000$ simulation). Notations as in Figure 3. Best periods (diamonds) are explained in Section 3.7.
@@ 3.6 Model 6
We conclude that the DFT “detects” the $P_{1}/2$ and $P_{1}$ periods, where $P_{1}$ is the correct simulated period value. However, the DFT two signal model is not the correct model for these Model 6 simulated data, which contain only one signal. If the correct period is $P$ and the correct model is a double wave $(K_{2}=2)$, the DCM pure sine model $(K_{1}=1)$ analysis can also give the values $P/2$ and $P$.
@@ 3.7 Model 7
Table 8: Fisher tests. DCM analysis of Model 1 simulated data ($n=100,{\mathrm{SN}}=100$, data file Model1n100SN100). (1) Simple model: Model number $\mathcal{M}$, model $g_{K_{1},K_{2},K_{3}}$, number of free parameters $\eta$, chi-square $\chi^{2}$ and control file name. (2-8) Complex model: Hypothesis H0 rejected $\uparrow$ (Complex model better), Hypothesis H0 not rejected $\leftarrow$ (Simple model better), test statistic $F$ and critical level $Q_{\mathrm{F}}$.22footnotetext: Note: No Fisher-test between models $\mathcal{M}$=4 and $\mathcal{M}$=5  because $\eta_{1}=\eta_{2}=6$.
@@ 3.8 Best model
Table 9: Fisher tests for Model 1 $n=100,{\mathrm{SN}}=100$ simulated data (electronic data file Model1n100SN100). (1) Polynomial model. (2) DCM model $g_{1,1,0}$. Note that $\mathcal{M}$=1-3  polynomials represent simple models, and $\mathcal{M}$=5-8  polynomials represent complex models. Otherwise as in Table 8. 33footnotetext: Note: No Fisher-test for $\mathcal{M}$=4  $(\eta_{1}=\eta_{2}=4)$.
@@
In the above example, the Fisher-test finds the correct number of signals $(K_{1})$ and the correct trend $(K_{3})$ for the pure sine signal alternative $(K_{2}=1)$. We do not test the double wave signal alternative $(K_{2}=2)$ against the pure sine signal alternative $(K_{1}=1)$ because the number of tested models would increase from 8 to 16, and Table 8 would become excessively large. One example of testing the $K_{2}=2$ signal models against each other can be found in Jetsu (2025, Table S7).
@@ 3.9 Significance estimates
There is no need to discuss the DFT significance estimates [17, their Equation 22] because this method fails to detect the correct Model 1 period.
@@ 3.10 Prediction
There are numerous techniques for predicting (forecasting) a time series [e.g., 16, 18]. The DCM model $g(t_{i},\bm{\beta})$ can be used to predict. We divide all data into the predictive data and the predicted data. The time points, the observations and the errors of these samples are
@@
during the chosen time interval $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$. Jetsu (2025) used this $m_{\mathrm{pred}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720. Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
@@ 4 Discussion
The solution for the non-linear DCM model is an ill-posed problem (Equations 1-5). We present a numerical solution that fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem. The DCM model is just one special case of a non-linear model. Our technique can be applied to solve other non-linear models: Use the free parameters that make the model non-linear (Equation 9: $\bm{\beta}_{I}$) for solving the remaining other free parameters (Equation 2.1: $\bm{\beta}_{II}$).
@@ 4.1 Existence ( $\mathrm{C}_{1}$)
[10] transformed the original function into the frequency domain. The modern DFT time series analysis method transforms the original data into the frequency domain and gives the best frequency for a pure sine model. [12] presented the least squares method, which minimises the differences between the data and the linear model. Our DCM does the same by testing a large number of linear models. The data spacing, even or uneven, is irrelevant for these models. For every chosen DCM model, the number of tested linear models is